OpenStudy (anonymous):

when the cost of a car is multiplied by 0.06 the result is $1780. find the cost of the car.

OpenStudy (anonymous):

x * 0.06 = 1780 x = 1780 / 0.06 x = 29667(rounded to the nearest whole number)

OpenStudy (anonymous):

Well, I have to explain it to you though. So, what is the inverse of multiplication? Division. So do long division with 1780 divided by 0.06= accurately, 296667